OpenStudy (anonymous):

A carpenteer is building a rectangle room with fixed perimeter of 80 feet, what dimensions would yeild the maximum area? What is the maximum area?

OpenStudy (anonymous):

w L = area Perimeter is: 2 w + 2 L = 80 L = 40 - w from the perimeter. w (40-w) = area Derivative of w (40-w) set to 0 is 40-2w = 0 w = 20 The maximum area is a 20 by 20 square of 400 square feet